.   If 35.0 g H2O at 22.7°C is combined with 65.0 g H2O at 87.5°C, what is the final temperature of the 

      mixture? The specific heat capacity of water is 4.184 J/g⋅°C.


1
Expert's answer
2022-04-21T07:45:04-0400

Solution:

This problem can be summarized thusly:

qgained by cold water = qlost by hot water

 

q = m × C × ΔT

q = m × C × (Tf − Ti),

where:

q = amount of heat energy gained or lost by substance (J)

m = mass of sample (g)

C = specific heat capacity (J °C−1 g−1)

Tf = final temperature (°C)

Ti = initial temperature (°C)


qgained by cold water = (35.0 g) × Cwater × (Tf − 22.7°C)

qlost by hot water = (65.0 g) × Cwater × (87.5°C − Tf)

Therefore,

(35.0) × Cwater × (Tf − 22.7) = (65.0) × Cwater × (87.5 − Tf)

(35.0) × (Tf − 22.7) = (65.0) × (87.5 − Tf)

35.0Tf − 794.5 = 5687.5 − 65.0Tf

100Tf = 6482

Tf = 64.82°C


Answer: The final temperature of the mixture is 64.82°C